Write the slope-intercept equation for the line that passes through (-1, -4) and (4, 11). Please show all of your work so that I may understand slopes.
Answer by rfer(12644)

slope m=(11--4)/(4--1)
m=15/5
m=3
-------------------
y=mx+b
11=3*4+b
11=12+b
b=11-12
b=-1
--------------
y=3x-1